OpenStudy (3m0k177y):

Evaluate 48 x a to the negative 3rd power for a = 2 8 6 One-Eighth One-Sixth

OpenStudy (coolcat123):

\[48\times48\times48\times2\]

563blackghost (563blackghost):

\(\huge\bf{48 \times 2^{-3}}\) Thats how it looks right?

563blackghost (563blackghost):

\(\huge\bf{a^{-b}=\frac{1}{a^{b}}}\) So we plug in.... \(\huge\bf{2^{-3}=\frac{1}{2^{3}}}\)

563blackghost (563blackghost):

So you would do.... \(\huge\bf{48 \times \frac{1}{2^{3}}}\)